### v2

# Maneuvering measurement-coherence into measurement-entanglement

Abstract
Quantum dynamics governs the transformation of static quantum resources, such as coherence and entanglement, in both quantum states and measurements. Prior studies have established that a quantum channel’s state-cohering power can be converted into state-entangling power without additional coherence. Here, we complete this coherence-to-entanglement paradigm by demonstrating that a channel’s measurement-cohering power can likewise be converted into measurement-entangling power. This result reinforces, on the dynamical level, the intuition that entanglement emerges as a manifestation of coherence. To formalize this picture, we develop resource theories for measurement-cohering and measurement-entangling powers and characterize the structure of incoherent measurements to analyze measurement-coherence generation. Furthermore, we show that the state-cohering power of a quantum channel is equivalent to the measurement-cohering power of its adjoint map, and a corresponding equivalence also exists between state-entangling power and measurement-entangling power.
Introduction.—Quantum coherence, deeply rooted in the linear structure of quantum theory [1], characterizes quantum superposition and underlies iconic phenomena such as Schrödinger’s cat and the double-slit interference. Its role extends broadly across diverse quantum technologies, including quantum sensing [2, 3], quantum computation [4, 5, 6], quantum correlations [7], and quantum thermodynamics [8, 9]. Quantum coherence has thus been rigorously formalized within quantum resource theory [10, 11, 12, 13, 14], paralleling the framework initially developed for entanglement. Quantum coherence is also a prerequisite for quantum entanglement, which can be intuitively understood as coherence distributed among multiple quantum subsystems. This fundamental connection has inspired extensive studies exploring the conversion of state-coherence into state-entanglement under incoherent operations within continuous-variable optical systems [15, 16, 17, 18] as well as finite-dimensional systems [19, 20, 21, 22].
This coherence-entanglement relationship extends beyond quantum states to quantum measurements. Quantum measurements themselves hold independent significance as carriers of quantum resources [23, 24, 25, 26, 27, 28, 29, 30]. Quantum coherence in measurements—measurement-coherence—is indispensable for comprehensively probing quantum states with coherence [27, 31, 32]. Similarly, quantum entanglement in measurements—measurement-entanglement—offers advantages in quantum information tasks such as quantum state discrimination and quantum system learning [28, 33, 34]. Recently, measurement-coherence has also been shown capable of converting into measurement-entanglement [35].
Figure 1: Quantum dynamics can modify static quantum resources either in quantum states or in quantum measurements.
Quantum dynamics, equally fundamental in quantum theory alongside states and measurements, plays a more intricate role concerning quantum resources, as depicted in Fig. 1. Quantum dynamics can create [36, 37, 38], preserve [39], or destroy static quantum resources [40]. Beyond merely altering resources, dynamics themselves constitute independent quantum resources capable of simulating other processes that lack intrinsic quantum resources [41, 42]. While the interplay between quantum dynamics and static quantum resources significantly enriches resource theories, notably, it has recently been demonstrated that certain aspects of quantum dynamics related to static quantum resources parallel the relationship between static coherence and entanglement. Specifically, it has been shown that the state-cohering power of quantum dynamics can be converted into state-entangling power [43].
In this Letter, we complete this foundational coherence-entanglement relationship at the dynamical level by introducing and analyzing a resource theory for quantum channels based on their ability to generate measurement-resources, specifically measurement-cohering power and measurement-entangling power. We establish a fundamental relation showing that the measurement-cohering power of quantum dynamics can be converted into measurement-entangling power. To achieve these results, we demonstrate that all incoherent measurements reduce to canonical incoherent basis measurements. Furthermore, we reveal an equivalence between the state-cohering power of a unital quantum channel and the measurement-cohering power of its adjoint channel, accompanied by a corresponding equivalence for entanglement. Our findings thus provide a unified and consistent picture of how quantum dynamics’ capacity to generate static coherence can be converted into the capacity to generate static entanglement, static resources in the form of either state-resources or measurement-resources, thereby significantly broadening the conceptual scope of quantum resource theories.
Measurement-coherence and measurement-entanglement.—In the resource theory of quantum coherence [14], quantum coherence of an operator is quantified with respect to a fixed orthonormal basis, which is referred to as the incoherent basis. The choice of an incoherent basis is contingent upon the experimental or physical constraints imposed on the system under consideration. We set an incoherent basis of a system $A$ and denote it as $\mathcal{I}_{A}=\{\lvert i\rangle_{A}:i=0,\dots,d-1\}$, with respect to which quantum coherence is considered. An operator $O_{A}$ that is diagonal in the incoherent basis is referred to as incoherent. Any incoherent operator is a fixed point under the dephasing channel $\Delta_{A}(O_{A})=\sum_{i=0}^{d-1}\langle i\rvert O_{A}\lvert i\rangle_{A}\left|i\right\rangle\!\left\langle i\right|_{A}=O_{A}$. Similarly, a quantum state $\rho_{A}$ is incoherent if $\Delta_{A}(\rho_{A})=\rho_{A}$. In this Letter, we consider constituent systems of dimension $d$ and assume that a quantum measurement for each system has $n$ outcomes, which can be described as a positive operator valued measure (POVM) given by $\mathcal{M}_{A}=\{M_{x}:\sum_{x=0}^{n-1}M_{x}=I_{A},M_{x}\geq 0\;\forall x=0,\ldots,n-1\}$, where $I_{A}$ is the identity operator. A quantum measurement $\mathcal{M}_{A}$ can be equivalently described by a quantum channel with classical outputs as
$\mathcal{M}_{A}(\rho_{A})=\sum_{x=0}^{n-1}\operatorname{Tr}_{A}(M_{x}\rho_{A})\left|x\right\rangle\!\left\langle x\right|_{A^{\prime}},$ (1)
where $\rho_{A}$ is a quantum state and $A^{\prime}$ denotes the output system [44]. The output of a quantum measurement is regarded as being classical implying that $\{\left|x\right\rangle\!\left\langle x\right|_{A^{\prime}}:x=0,\dots,n-1\}$ is an incoherent basis for the system $A^{\prime}$. A quantum measurement $\mathcal{M}_{A}$ is called an incoherent measurement if its POVM elements are incoherent, that is, $\Delta_{A}(M_{x})=M_{x}$ for all $x=0,\dots,n-1$. If a quantum measurement $\mathcal{M}_{A}$ is incoherent, the probability of an outcome is solely dependent on the incoherent information of input states, disregarding the coherent components [27, 45], as demonstrated by the following equation:
$p(x)=\operatorname{Tr}_{A}M_{x}\rho_{A}=\operatorname{Tr}M_{x}\Delta_{A}(\rho_{A}).$ (2)
We denote the set of incoherent measurements as $\mathbf{IM}$. Measurement-coherence of a quantum measurement $\mathcal{M}_{A}$ is quantified by the relative entropy of measurement-coherence $C_{m}(\cdot)$, as defined in [35]:
$C_{m}(\mathcal{M}_{A})=\min_{\mathcal{F}_{A}\in\mathbf{IM}}D_{m}(\mathcal{M}_{A}\|\mathcal{F}_{A}),$ (3)
where the measurement relative entropy between two quantum measurements $D_{m}(\cdot\|\cdot)$ is proportional to the average relative entropy between POVM elements as
$D_{m}(\mathcal{M}_{A}\|\mathcal{N}_{A})\coloneqq\dfrac{1}{d}\sum_{x=0}^{n-1}D(M_{x}\|N_{x}),$ (4)
with the quantum relative entropy $D(\cdot\|\cdot)$. The relative entropy of measurement-coherence of a quantum measurement is found to be equal to the average relative entropy of coherence of each POVM element, up to a multiplicative constant [35]:
$C_{m}(\mathcal{M}_{A})=\dfrac{1}{d}\sum_{x=0}^{n-1}\{S(\Delta_{A}(M_{x}))-S(M_{x})\},$ (5)
where $S(X_{A})=-\operatorname{Tr}_{A}X_{A}\log X_{A}$ for $X_{A}\geq 0$ is the von Neumann entropy.
Figure 2: A quantum measurement $\mathcal{M}_{A}$ transforms to another quantum measurement ~$\widetilde{\mathcal{M}}_{A}$ under a pre-processing channel $\mathcal{E}_{A}$ and a classical post-processing channel $\mathcal{S}_{A^{\prime}}$. The pre-processing channel can generate measurement-resource.
A quantum measurement $\mathcal{M}_{A}$ undergoes a transformation into another quantum measurement ~$\widetilde{\mathcal{M}}_{A}$ when a pre-processing channel and a classical post-processing channel are added, as illustrated in Fig. 2 [23]. This can be described as follows:
~$\widetilde{\mathcal{M}}_{A}=\mathcal{S}_{A^{\prime}}\circ\mathcal{M}_{A}\circ\mathcal{E}_{A},$ (6)
where a classical post-processing channel $\mathcal{S}_{A^{\prime}}(\rho_{A^{\prime}})=\sum_{x,x^{\prime}=0}^{n-1}p(x^{\prime}|x)\langle x\rvert\rho_{A^{\prime}}\lvert x\rangle_{A^{\prime}}\left|x^{\prime}\right\rangle\!\left\langle x^{\prime}\right|_{A^{\prime}}$ sends an outcome $x$ to another outcome $x^{\prime}$ with a conditional probability distribution $p(x^{\prime}|x)$ satisfying $\sum_{x^{\prime}}p(x^{\prime}|x)=1$ for each $x$. At this point, we demonstrate that an incoherent measurement is, in essence, the incoherent basis measurement:
Theorem 1.
An incoherent measurement $\mathcal{M}_{A}\in\mathbf{IM}$ can be implemented by the incoherent basis measurement $\mathcal{I}_{A}=\{\left|i\right\rangle\!\left\langle i\right|_{A}:i=0,\dots,d-1\}$ followed by a classical post-processing channel $\mathcal{S}_{A}$:
$\mathcal{M}_{A}=\mathcal{S}_{A}\circ\mathcal{I}_{A}.$ (7)
We defer proofs of the above and subsequent results to the Supplemental Material. The set of quantum channels that do not generate any measurement-coherence is known to be detection-incoherent operations ($\mathbf{DIO}$) [27]. These serve as free operations for measurement-coherence. A quantum channel $\mathcal{E}_{A}$ is in $\mathbf{DIO}$ if and only if it satisfies the equation $\Delta_{A}\circ\mathcal{E}_{A}=\Delta_{A}\circ\mathcal{E}_{A}\circ\Delta_{A}$. This characterization is analogous to the characterization of a quantum channel in the maximal incoherent operations ($\mathbf{MIO}$), which is the largest set of quantum channels that send incoherent states to incoherent states.
In the resource theory of measurement-entanglement, we take the closure of the set of LOCC measurements ($\overline{\mathbf{LOCCM}}$) as our free resources, that can be implemented as LOCC operations with local measurements [44]. The closure of the set of LOCC channels ($\overline{\mathbf{LOCC}}$) corresponds to the free operations that preserve $\overline{\mathbf{LOCCM}}$. Measurement-entanglement of a bipartite quantum measurement $\mathcal{M}_{AB}$ is quantified by the relative entropy of measurement-entanglement of $\mathcal{M}_{AB}$ defined as follows  [^1] We choose $\overline{\mathbf{LOCCM}}$ as our free resource, which possesses a more direct operational meaning than $\mathbf{SepM}$, the free resource used in [35].:
$E_{m}(\mathcal{M}_{AB})\coloneqq\min_{\mathcal{F}\in\overline{\mathbf{LOCCM}}(A:B)}D_{m}(\mathcal{M}_{AB}\|\mathcal{F}_{AB}).$ (8)
We remark that, despite the similarities between the resource theories of quantum coherence and entanglement, there is no entanglement-destroying channel that destroys entanglement while holding every separable state as its fixed point [47]. This is in contrast to the case of quantum coherence, where the dephasing channel destroys static quantum coherence while holding every incoherent state as a fixed point of the channel. As a result of this limitation, there is no straightforward operational interpretation or concise formula for $E_{m}$ as there is for the relative entropy of measurement-coherence reduced in Eq. (5).
Measurement-resource powers of quantum dynamics.—Now we focus on the capacity of quantum dynamics to generate measurement-resources. Quantum channels are capable of generating measurement-coherence and measurement-entanglement by acting as a pre-processing channel prior to a quantum measurement. In contrast, classical post-processing after a quantum measurement is unable to generate any measurement-resources.
As previously introduced, $\mathbf{DIO}$ is the largest set of quantum channels under which incoherent measurements remain incoherent [35]. In light of this, we define the measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ as follows:
$C(\mathcal{E}_{A})\coloneqq\min_{\mathcal{F}_{A}\in\mathbf{DIO}}\max_{\mathcal{M}_{A\widetilde{A}}\in\mathbf{IM}}D_{m}(\mathcal{M}_{A\widetilde{A}}\circ\mathcal{E}_{A}\|\mathcal{M}_{A\widetilde{A}}\circ\mathcal{F}_{A}),$ (9)
where the ancillary system $\widetilde{A}$ can be arbitrary. In fact, we find that the measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ is equal to the amount of the measurement-coherence generated by the quantum channel $\mathcal{E}_{A}$ from the incoherent basis measurement $\mathcal{I}_{A}$:
Theorem 2.
For a quantum channel $\mathcal{E}_{A}$, the measurement-cohering power of the quantum channel is given by
$C(\mathcal{E}_{A})=C_{m}(\mathcal{I}_{A}\circ\mathcal{E}_{A}),$ (10)
where $\mathcal{I}_{A}=\{\left|i\right\rangle\!\left\langle i\right|_{A}:i=0,\dots,d-1\}$ is the incoherent-basis measurement.
Similarly, the free resource in measurement-entangling power corresponds to $\overline{\mathbf{LOCC}}$ channels. We define the measurement-entangling power of a quantum channel $\mathcal{E}_{AB}$ as follows:
$E(\mathcal{E}_{AB})\coloneqq\min_{\mathcal{F}\in\overline{\mathbf{LOCC}}(A:B)}\max_{\mathcal{M}\in\overline{\mathbf{LOCCM}}(A\widetilde{A}:B\widetilde{B})}\\ D_{m}(\mathcal{M}_{A\widetilde{A}B\widetilde{B}}\circ\mathcal{E}_{AB}\|\mathcal{M}_{A\widetilde{A}B\widetilde{B}}\circ\mathcal{F}_{AB}),$ (11)
where the ancillary systems $\widetilde{A}$ and $\widetilde{B}$ can be arbitrary. As previously noted, there is no resource-destroying dynamics that fixes any separable state in entanglement theory. Thus, the aforementioned simplification as in Eq. (10) is not applicable.
The measurement-resource powers in Eq. (9) and Eq. (11) are legitimate resource monotones that satisfy the following properties [48]: i) the measurement-cohering power $C(\mathcal{E}_{A})$ is non-negative and faithful, that is, $C(\mathcal{E}_{A})$ is zero if and only if $\mathcal{E}_{A}$ is in $\mathbf{DIO}$. ii) The measurement-cohering power is monotone under a $\mathbf{DIO}$ $\mathcal{K}_{A}$ and a unital $\mathbf{DIO}$ $\mathcal{L}_{A}$ such that $C(\mathcal{K}_{A}\circ\mathcal{E}_{A}\circ\mathcal{L}_{A})\leq C(\mathcal{E}_{A})$. iii) The measurement-cohering power is convex, that is, it holds that $C(p\mathcal{E}_{A}+(1-p)\mathcal{G}_{A})\leq pC(\mathcal{E}_{A})+(1-p)C(\mathcal{G}_{A})$ for quantum channels $\mathcal{E}_{A}$ and $\mathcal{G}_{A}$, and any $0\leq p\leq 1$.
Similarly, the measurement-entangling power $E(\mathcal{E}_{AB})$ also satisfies the three properties: i) $E(\mathcal{E}_{AB})\geq 0$, with equality holding if and only if $\mathcal{E}_{AB}\in\overline{\mathbf{LOCC}}(A\!:\!B)$. ii) The measurement-entangling power $E(\mathcal{E}_{AB})$ is monotone under the composition by a $\overline{\mathbf{LOCC}}$ channel $\mathcal{K}_{AB}$ and a unital $\overline{\mathbf{LOCC}}$ channel $\mathcal{L}_{AB}$ such that $E(\mathcal{K}_{AB}\circ\mathcal{E}_{AB}\circ\mathcal{L}_{AB})\leq E(\mathcal{E}_{AB})$. iii) The measurement-entangling power $E$ is convex, that is, for quantum channels $\mathcal{E}_{AB}$ and $\mathcal{G}_{AB}$, and any $0\leq p\leq 1$, it holds that $E(p\mathcal{E}_{AB}+(1-p)\mathcal{G}_{AB})\leq pE(\mathcal{E}_{AB})+(1-p)E(\mathcal{G}_{AB})$.
Conversion of measurement-cohering power to measurement-entangling power.—Having established the concepts of measurement-cohering and measurement-entangling powers of quantum channels, we now present our main result: quantum coherence can be converted to entanglement at the level of dynamics generating measurement resources as illustrated in Fig. 3(a). First, we demonstrate that the measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ serves as an upper bound on the measurement-entangling power of a composite channel constructed from the quantum channel $\mathcal{E}_{A}$, without any additional measurement-cohering power:
Theorem 3.
Let $\mathcal{E}_{A}$ be a quantum channel. Then we have that
$E(\Delta_{AB}\circ\mathcal{K}_{AB}\circ\mathcal{E}_{A}\circ\mathcal{L}_{AB})\leq C(\mathcal{E}_{A})$ (12)
for a $\mathbf{DIO}$ channel $\mathcal{K}_{AB}$ and a unital $\mathbf{DIO}$ channel $\mathcal{L}_{AB}$.
Given that the quantum channels surrounding the quantum channel $\mathcal{E}_{A}$ are $\mathbf{DIO}$s, they do not contribute measurement-cohering power (See Fig. 3(a)). The dephasing channels in the post-processing serve to render the subsequent measurement incoherent and align the incoherent basis, thereby enabling the foregoing channels to convert the measurement-cohering power of the quantum channel $\mathcal{E}_{A}$ into the measurement-entangling power in that incoherent basis. In contrast, when converting state-cohering power to state-entangling power, dephasing channels serve as a preprocessing step to ensure the input states are incoherent, as depicted in Fig. 3(b) [43]. Consequently, Theorem 3 implies that measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ is the sole source of measurement-entangling power of the composite quantum channel. Indeed, there exists a configuration that allows for complete conversion of measurement-cohering power to measurement-entangling power although we restrict the pre-processing channel $\mathcal{L}_{AB}$ to be unital, as the following result shows (see Table 1):
Theorem 4.
For a quantum channel $\mathcal{E}_{A}$, its measurement-cohering power can be fully converted to measurement-entangling power under a pre-processing by the adjoint channel of the generalized CNOT channel $\mathcal{U}_{\mathrm{CNOT}}^{{\dagger}}$ and a post-processing dephasing channel:
$E(\Delta_{AB}\circ\mathcal{E}_{A}\circ\mathcal{U}_{\mathrm{CNOT}}^{{\dagger}})=C(\mathcal{E}_{A}),$ (13)
where $U_{\mathrm{CNOT}}=\sum_{i,j=0}^{d-1}\left|i,i\oplus j\right\rangle\!\left\langle ij\right|_{AB}$ with $\oplus$ denoting addition modulo $d$.
Figure 3: (a) Measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ can convert to measurement-entangling power of a quantum channel ~$\widetilde{\mathcal{E}}_{AB}$ under a $\mathbf{DIO}$ channel $\mathcal{K}_{AB}$, a unital $\mathbf{DIO}$ channel $\mathcal{L}_{AB}$, and the dephasing channel $\Delta_{AB}$. (b) State-cohering power of a quantum channel $\mathcal{E}_{A}$ can convert to state-entangling power of a quantum channel ~$\widetilde{\mathcal{E}}_{AB}$ under $\mathbf{MIO}$ channels $\mathcal{K}_{AB}$ and $\mathcal{L}_{AB}$, and the dephasing channel $\Delta_{AB}$ [43].
This result completes the picture of quantum coherence convertible to quantum entanglement at the dynamic level, specifically by addressing the previously unexplored role of quantum channels on the half of the static resources: measurement-resources. Measurement resources constitute half of the static resources in quantum information—complementing the widely studied state resources—yet have received comparatively less attention. This critical balance between measurement and state resources is fundamental, as state-coherence itself is inherently incomplete without coherent measurement to observe its effects [27, 32]. Similarly, measurement-entanglement is crucial in achieving exponential quantum advantage in quantum learning [33, 49, 50, 34].
The convertibility of measurement-cohering power to measurement-entangling power implies that a measurement-entangling power monotone induces a measurement-cohering power monotone, as demonstrated by the following result:
Theorem 5.
A measurement-entangling power monotone $\widetilde{E}$ induces a measurement-cohering power monotone $\widetilde{C}$ given by
$\widetilde{C}(\mathcal{E}_{A})=\max_{\mathcal{K}_{AB},\mathcal{L}_{AB}}\widetilde{E}(\Delta_{AB}\circ\mathcal{K}_{AB}\circ\mathcal{E}_{A}\circ\mathcal{L}_{AB}),$ (14)
where the maximization is over $\mathbf{DIO}$ channels $\mathcal{K}_{AB}$ and unital $\mathbf{DIO}$ channels $\mathcal{L}_{AB}$.
Our results reproduce those of static measurement-resources when considering measurement-resource powers for a quantum measurement described as a quantum channel, as in Eq. (1). A quantum measurement $\mathcal{N}_{A}$ is a quantum channel whose output is incoherent, satisfying that $\mathcal{N}_{A}=\Delta_{A}\circ\mathcal{N}_{A}$. This, along with the observation that the dephasing channel corresponds to the incoherent basis measurement, leads to the following result:
Theorem 6.
For quantum measurements, the measurement-resource powers reduce to the relative entropies of measurement-resources: for a quantum measurement $\mathcal{N}_{A}$ and a bipartite quantum measurement $\mathcal{N}_{AB}$,
$C(\mathcal{N}_{A})=C_{m}(\mathcal{N}_{A}),\quad E(\mathcal{N}_{AB})=E_{m}(\mathcal{N}_{AB}).$ (15)
We now proceed to investigate the relation between the resource powers of quantum channels with respect to both static quantum resources.
State-resource power versus measurement-resource power.—A quantum channel $\mathcal{E}_{A}$ can generate static quantum resources in both quantum states and measurements. The state-cohering power of $\mathcal{E}_{A}$, quantifying the maximum coherence the channel can induce in incoherent states, is given by
$C_{g}(\mathcal{E}_{A})\coloneqq\max_{\sigma_{A\widetilde{A}}\in\mathbf{I}}C_{R}(\mathcal{E}_{A}(\sigma_{A\widetilde{A}})),$ (16)
where $C_{R}(\rho_{A})\coloneqq\min_{\sigma_{A}\in\mathbf{I}}D(\rho_{A}\|\sigma_{A})$ is the relative entropy of coherence [12], and $\mathbf{I}$ denotes the set of incoherent states. The measurement-cohering power of a channel $\mathcal{E}_{A}$ is given in Eq. (9).
Are there relations between the two static resource powers of a quantum channel? In general, there is no relation between a channel’s state-resource power and its measurement-resource power. Consider a preparation channel $\mathcal{E}_{A}(\rho_{A})=\operatorname{Tr}_{A}(\rho_{A})\left|+\right\rangle\!\left\langle+\right|_{A}$. This quantum channel has the maximum state-cohering power as $C_{g}(\mathcal{E}_{A})=1$ but no measurement-cohering power as $C(\mathcal{E}_{A})=0$. Another quantum channel $\mathcal{G}_{A}(\rho_{A})=\langle+\rvert\rho_{A}\lvert+\rangle_{A}\left|0\right\rangle\!\left\langle 0\right|_{A}+\langle-\rvert\rho_{A}\lvert-\rangle_{A}\left|1\right\rangle\!\left\langle 1\right|_{A}$ shows the opposite case where it can generate no state-coherence as $C_{g}(\mathcal{G}_{A})=0$ while it can generate the maximum measurement-coherence $C(\mathcal{G}_{A})=1$. However, for a unital quantum channel, we establish a connection between state-cohering power and measurement-cohering power:
Theorem 7.
For a unital quantum channel $\mathcal{E}_{A}$, the state-cohering power $C_{g}(\mathcal{E}_{A})$ and the measurement-cohering power $C(\mathcal{E}_{A}^{{\dagger}})$ of its adjoint channel $\mathcal{E}_{A}^{{\dagger}}$ satisfy the following:
$\dfrac{1}{d}C_{g}(\mathcal{E}_{A})\leq C(\mathcal{E}_{A}^{{\dagger}})\leq C_{g}(\mathcal{E}_{A}).$ (17)
This result implies that for a unital quantum channel, the state-cohering power of the channel and the measurement-cohering power of its adjoint channel are equivalent in the sense that one is non-zero if and only if the other is non-zero.
For entanglement, we find a similar relation. The state-entangling power of a quantum channel $\mathcal{E}_{AB}$  [^2] The state-entangling power has also been used to characterize the average state-entanglement generated from arbitrary product states as in [36]. We follow the convention of [41] for consistency with the state-cohering power. is defined as
$E_{g}(\mathcal{E}_{AB})\coloneqq\max_{\sigma\in\mathbf{SepD}(A\widetilde{A}:B\widetilde{B})}E_{R}(\mathcal{E}_{AB}(\sigma_{A\widetilde{A}B\widetilde{B}})),$ (18)
where $E_{R}(\rho_{AB})\coloneqq\min_{\sigma\in\mathbf{SepD}(A:B)}D(\rho_{AB}\|\sigma_{AB})$ is the relative entropy of entanglement [52]. The state-entangling power of a unital quantum channel and the measurement-entangling power of its adjoint channel are equivalent in the following sense:
Theorem 8.
For a unital quantum channel $\mathcal{E}_{AB}$, the state-entangling power $E_{g}(\mathcal{E}_{AB})$ and the measurement-entangling power $E(\mathcal{E}_{AB}^{{\dagger}})$ of its adjoint channel $\mathcal{E}_{AB}^{{\dagger}}$ satisfy the following:
$\dfrac{1}{d^{4}}E_{g}(\mathcal{E}_{AB})\leq E(\mathcal{E}_{AB}^{{\dagger}})\leq E_{g}(\mathcal{E}_{AB}).$ (19)
Conclusion.—We have established resource theories quantifying the measurement-cohering and measurement-entangling powers of quantum channels, demonstrating a conversion between these two measurement-related dynamical resources. This result completes the conceptual framework linking quantum coherence and quantum entanglement at both static and dynamic levels, as illustrated in Table 1. Specifically, while previous studies addressed the connection between quantum channel’s state-resource powers, our work connects the measurement-cohering power of a quantum channel to the measurement-entangling power. Our results integrate and extend prior studies on the measurement resources by treating quantum measurements as channels producing classical data. This approach reproduces known results on conversion from measurement-coherence to measurement-entanglement, placing them within our broader theoretical framework. In achieving our main results, we utilized the newly characterized structure of an incoherent measurement as an incoherent basis measurement followed by a classical post-processing channel.
State-resource Measurement-resource
Static [Uncaptioned image] [Uncaptioned image]
Dynamic [Uncaptioned image] [Uncaptioned image]
Table 1: Illustration of the optimal conversion of quantum coherence into entanglement at different levels of resource theory. The table categorizes these conversions into static (quantum states and measurements) and dynamic (quantum channels) scenarios. The diagrams, arranged from the top left to the bottom left in a clockwise direction, depict: the optimal conversion of quantum coherence of a quantum state $\rho_{A}$ to entanglement of a quantum state $\widetilde{\rho}_{AB}$ [19], the optimal conversion of measurement-coherence of a quantum measurement $\mathcal{M}_{A}$ to measurement-entanglement of a quantum measurement ~$\widetilde{\mathcal{M}}_{AB}$ [35], the optimal conversion of measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ to measurement-entangling power of a quantum channel ~$\widetilde{\mathcal{E}}_{AB}$ (this work), and the optimal conversion of state-cohering power of a quantum channel $\mathcal{E}_{A}$ to state-entangling power of a quantum channel ~$\widetilde{\mathcal{E}}_{AB}$ [43].
Moreover, we demonstrated that the state-cohering power of a unital quantum channel is equivalent to the measurement-cohering power of its adjoint channel, and that a similar equivalence holds between the state-entangling power of a unital bipartite quantum channel and the measurement-entangling power of its adjoint channel.
Despite significant advances, fundamental open questions persist. While coherence-to-entanglement conversion is well-documented, the reverse transformation remains poorly understood due to the basis-independent nature of entanglement, making such conversions inherently nontrivial. Moreover, the interplay between quantum coherence and multipartite entanglement remains largely unexplored [21], partly owing to the incomplete theoretical characterization of multipartite entanglement itself.
Another critical point is that static resource generation alone cannot capture the full complexity of quantum dynamics concerning resource utilization. Indeed, quantum dynamics itself can be viewed as a resource under appropriate restrictions on quantum superchannels [53, 54, 55, 56]. For example, the generalized Hadamard gate has been investigated as a dynamic coherence resource [57], while the SWAP gate functions as a dynamic entanglement resource capable of simulating arbitrary bipartite quantum channels [42, 56, 58], analogous to how Bell states serve as universal resources for state preparation. This viewpoint connects naturally to quantum simulation frameworks [59, 60, 61], one of quantum computing’s most promising applications, emphasizing the essential relationship between experimental capabilities and required quantum resources. However, the connections among these dynamic quantum resources remain less understood compared to their static counterparts. A deeper insight into these dynamical relationships holds potential for developing more efficient algorithms, particularly in quantum simulation.
Addressing these open issues will significantly enhance our fundamental understanding of quantum resources and promote the advancement of resource-efficient quantum information processing technologies.
Acknowledgements.
H-.J. K acknowledges was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (2023M3K5A1094813). S. L. acknowledges support from the National Research Foundation of Korea (NRF) Grants funded by the Ministry of Science and ICT (MSIT) (RS-2022-NR068791, RS-2024-00432214, and RS-2022-NR073109), and Creation of the Quantum Information Science R&D Ecosystem (RS-2023-NR068116) through the NRF funded by the MSIT.